OpenStudy (anonymous):

Idk how to complete the square. :( Trying to get x^2-16x+56=y into VERTEX FORM. So far i have x^2-16x=-56 ... Help? :/

OpenStudy (anonymous):

add 64 on both sides

OpenStudy (anonymous):

Where did you get 64? :/

OpenStudy (anonymous):

-16/2 = -8 (-8)^2 = 64

OpenStudy (anonymous):

okay. Now i have x^2-16x+64=8 I'm sorry, this is annoying but i don't know what to do next.

OpenStudy (anonymous):

(x - 8)^2 = 8

OpenStudy (anonymous):

Thank you. (x-8)^2=8 is vertex form?

OpenStudy (anonymous):

(x-8)^2 - 8 = 0